# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ 3.3.2 Non-imaging-based feature learning
Most non-imaging data is presented in the form of text. Before fusion with image data, text data needs to be embedded. We first introduce a widely-used basic text embedding technique, the lookup table. This table assigns a unique index to each word or character, which is used to look up a pre-trained word vector or character vector. In addition to the basic method, additional feature extraction can be performed on these vectors to enhance their representation capabilities.
Transformer-based models such as BERT and its variants have emerged as mainstream methods for feature extraction across various textual data, such as terminology (Liu et al., 2021e; Cao et al., 2022, 2023; Liu et al., 2021b; Xue et al., 2024; Liu et al., 2023b; Li et al., 2023c), real text reports (Liu et al., 2021b, 2023b; Li et al., 2023c; Jin et al., 2024), knowledge graphs (Yang et al., 2022; Huang et al., 2023; Li et al., 2023a; Xu et al., 2023), and questionnaires (Tanwani et al., 2022; Pellegrini et al., 2023), and have achieved good results.
Several methods are designed for a specific type of input. Li et al. (2023c) used the TF-IDF to re-weight terminology embeddings. The re-weighted approach alleviated the issue of data imbalance, resulting in performance enhancement. Clinical information can be processed by a pre-trained feature extractor named BioSentVec (Zhou et al., 2021; Chen et al., 2019). For the knowledge base, Yang et al. (2022) used a knowledge graph embedding model named RotatE (Sun et al., 2019b) to obtain entity embeddings and relation embeddings from the RadGraph. Besides using the entire RadGraph as input, two works combined the real reports with RadGraph to extract case-related information from real reports and queried related information from the RadGraph (Yang et al., 2022; Li et al., 2023a). Alternatively, Li et al. (2023c) considered RadGraph as an annotation tool for extracting entities and positional information from real reports. Other two works (Xu et al., 2023; Huang et al., 2023) utilized classification results to process the self-built graph and extracted case-related information. The experimental results substantiate the beneficial impact of incorporating case-related knowledge on report generation, particularly evident when integrating real reports with the knowledge base. In addition, age and gender information are not text data and are generally encoded as one-hot vector (Zhou et al., 2021).
@@ 3.4 Multi-modal feature fusion and interaction
Feature fusion and interaction refer to the integration of multi-modal data from inputs or auxiliary tasks. This step has two purposes. First, visual and textual features from disparate domains present challenges for model learning. By fusing and facilitating interaction between these features, the domain gap can be narrowed, thereby enhancing network learning. Second, the image regions should align with the sentences in the reports. This correspondence can be learned through fusion and interaction. In the auxiliary task of embedding comparison (see Section 3.3.1), semantic features extracted from real reports are used to supervise the learning of image features. However, this approach differs from multi-modal feature fusion. The objective of embedding comparison is to enhance image features, without incorporating non-image features into the generator. Instead, the fused features in this section are forwarded to the generator.
The most straightforward approach for feature fusion and interaction is feature-level operation including the concatenation, summation, or multiplication of multimodal features (13 works). However, the feature-level operation could be too simple to enable sufficient interaction. Therefore, neural network-based methods are leveraged, such as LSTMs (2 works) and the multi-head attention mechanism (27 works). While this approach facilitates convenient feature fusion, its lack of specific design for multi-modality data fusion leads to limited effectiveness in interaction.
@@
We would like to highlight a memory metric-based method proposed by Chen et al. (2021). It significantly enhanced the performance of the report generation system by facilitating feature interaction. A metric was initialized randomly. Image features, text features from generated tokens, and memory metric features were then projected into the same space. Subsequently, distances between image features and memory metric features, as well as text features and memory metric features, were calculated. The top K metric features with the closest distances to the image or text were selected, respectively. These selected features were then weighted based on these distances and were fed back into an encoder-decoder structure. Two studies (Qin and Song, 2022; You et al., 2022) followed this method. Wang et al. (2022a) modified it in two ways: 1) they initialized the matrix by visual and textual features; 2) the cross-modal interaction occurred only among cases with the same label. These two modifications both resulted in a notable improvement. Li et al. (2023d) contended that the methods mentioned lack explicit constraints for cross-modal alignments. They considered orthonormal bases as the metric and input them along with visual or textual features, into multi-head attention modules. Then the outputs of attention modules were processed by a self-defined gate mechanism. A triplet matching loss was utilized to align the processed visual and textual features. This method slightly improved the results.
@@ 3.5 Report generation
The last step is report generation, which utilizes extracted features from earlier steps to produce the final reports. The generation methods mainly include decoder-based techniques (Section 3.5.1), retrieval-based techniques (Section 3.5.2), and template-based techniques (Section 3.5.2). In addition, the development of large language models has made it possible to utilize them to enhance the quality of generated reports. It is discussed in Section 3.5.3
3.5.1 Decoder-based techniques
The decoder decodes the extracted representation of inputs and generates a descriptive report. The mainstream architectures include LSTM (Hochreiter and Schmidhuber, 1997) and Transformer. Compared to LSTM, the Transformer processes the entire sequence simultaneously rather than sequentially. Therefore, the Transformer allows for more efficient parallelization during training and can capture long-range dependencies. In the 89 reviewed papers, the Transformer tends to replace LSTM. Fifty-five works utilized the Transformer as a decoder and only 23 of them utilized LSTMs (12 works) or hierarchical LSTMs (11 works). For the papers published in 2023 and 2024, all encoder-decoder structures used the Transformer as their decoders. There are two ways to modify the decoder and improve model performance.
@@ 3.5.1 Decoder-based techniques
Memory-driven Transformer: We would like to highlight the Memory-driven Transformer (R2Gen) proposed by Chen et al. (2020). It has been increasingly popular in recent years. The R2Gen introduces a memory module and a memory-driven conditional layer normalization module into the Transformer decoder architecture. The design of the memory module hypothesizes that diverse images exhibit similar patterns in their radiological reports, thereby serving as valuable references for each other. Building a memory matrix can capture this pattern and transfer it during training. Specifically, similar to that in the section 3.3.1, a matrix is randomly initialized and is updated using the gate mechanism based on the matrix from the last step and generated reports. The layer normalization is designed to integrate the outputs of the memory module into the decoder.
@@
3.5.2 Retrieval-based and template-based techniques

### v2

@@ 3.3.2 Non-imaging-based feature learning
The types of non-image data include text, graph, and numerical data (see Figure 3). Numerical data is typically fed into the network in the next step after one-hot encoding, without a specifically designed encoder. Therefore, in this section, we focus on feature extraction for text and graph inputs.
After tokenization mentioned in Section 3.2, text data is divided into basic units (such as words, subwords, or characters). Then the general embedding technique (e.g., lookup table) is applied to convert these tokens into continuous vector representations. Next, many studies directly fuse these vectors with image features, while others incorporate a text encoder to enhance the representational capabilities of textual features. Transformer-based models and their variants are mainstream methods, widely applied to terminology (Liu et al., 2021e; Cao et al., 2022, 2023; Liu et al., 2021b; Xue et al., 2024; Liu et al., 2023c; Li et al., 2023d), real text reports (Liu et al., 2021b, 2023c; Li et al., 2023d; Jin et al., 2024), and questionnaires (Tanwani et al., 2022; Pellegrini et al., 2023a). These models are typically trainable, and some studies also utilize frozen models that are pre-trained on medical datasets to extract clinical features more effectively. For example, Li et al. (2023d) utilized the Clinical Bidirectional Encoder Representations from Transformers (ClinicalBERT) (Alsentzer et al., 2019) to extract clinical concept features. There is no experimental evidence proving that the introduction of pre-trained models enhances performance; however, reducing the number of trainable parameters can accelerate convergence and facilitate training. In addition, in medical reports, there is an imbalance between normal and abnormal descriptions, which may cause the model to prioritize normal descriptions. To address this issue, Li et al. (2023d) employed TF-IDF to re-weight the input terminology embeddings. During training, the weights are derived from ground truth reports, while during inference, they are obtained from the reports of similar training images.
In knowledge graph feature extraction, extracting node embeddings with pre-trained transformer models is common (Huang et al., 2023; Li et al., 2023b; Xu et al., 2023); the critical challenge lies in effectively utilizing edge information (i.e., relationships). The pre-defined graph connecting organs and diseases indicates their association but does not specify the relationships. Traditional graph convolutional networks utilize fixed weights, such as the adjacency matrix, to represent the association and update the node features by averaging the features of neighboring nodes. However, this approach fails to capture the importance of different neighbors to the node. To address this, the graph attention mechanism is introduced and computes the attention weights between the node and its neighbors, then performs a weighted aggregation of the neighbors’ features to update the node’s feature (Huang et al., 2023; Xu et al., 2023; Li et al., 2023b). In addition, Li et al. (2023b) added identifiers to indicate node levels (i.e., root, organ, and disease). Combined with the case-specific extraction method (see Section 3.2), this approach yields significant performance improvements.
Yang et al. (2022) aims to represent more complex relationships between medical entities, utilizing a general graph embedding model called RotatE (Sun et al., 2019). Relationships are represented as rotations from entity1 to entity2 within a complex vector space. In this way, four common relationships (i.e., symmetry, antisymmetry, inversion, and composition) can be represented. However, these four relationships are not practical for medical data. Relationships can be more specified in the medical field; for instance, RadGraph defines relationships between medical entities as “suggestive of”, “located at”, and “modify”. Yang et al. (2022) transformed it into sentences (e.g., “pneumothorax suggestive of bleeding") to extract semantic features using a pre-trained transformer model.
In conclusion, feature extraction from text data typically relies on transformer models, which lack distinctive designs. For graph data, leveraging relationships is a crucial topic. Exploring ways to incorporate more specific relationships presents a promising research direction.
@@ 3.4 Multi-modal feature fusion and interaction
Feature fusion and interaction refer to the integration of multi-modal data from inputs or auxiliary tasks. It is important for report generation using multi-modal inputs. As discussed in Section 1, aligning multimodal data involves four main challenges: distinct feature spaces, identifying complementary and redundant information, various types of noise, and uneven contributions. Fusion and interaction methods can help alleviate the above problems. Additionally, the correspondence between image regions and report sentences can be learned during this process.
The most straightforward approach is feature-level operation including the concatenation, summation, or multiplication of multimodal features (14 works). However, the feature-level operation could be too simple to enable sufficient interaction. Therefore, neural network-based methods are leveraged. Huang et al. (2021a) used an LSTM network to integrate keywords and image features, starting with image features as input and sequentially introducing keywords at each time step. This method is relatively cumbersome and less effective than attention mechanisms. The multi-head attention mechanism is a widely adopted method for feature fusion (28 works), involving three inputs: query, key, and value. Its core idea is to compute the similarity between the query and key, and use this correlation to weight the value, enabling the identification of alignable features across different modalities while filtering out irrelevant information. The attention mechanism can be applied to concatenated feature data through self-attention, allowing for an arbitrary number of input modalities. For example, Dalla Serra et al. (2022) concatenated clinical information tokens with visual features and used the combined features as the query, key, and value. Pellegrini et al. (2023a) merged question features with image features, feeding the concatenated features into the BERT model. However, it does not differentiate between specific modalities. Alternatively, another approach is that one modality (e.g., images) can act as the query, while another modality (e.g., textual keywords) serves as keys and values, allowing the query modality to guide the model in learning to identify relevant information from other modalities. For instance, Yang et al. (2022) proposed a Knowledge-enhanced multi-head attention mechanism, using visual features as queries and knowledge graph entity embeddings as keys and values for feature fusion, with relation embeddings incorporated. Ablation studies showed that including relation embeddings improves performance. However, there is no detailed experimental analysis of the difference between concatenating features as Q, K, V inputs in attention and using one as Q and the others as K and V.
We would like to highlight the memory metric-based method proposed by Chen et al. (2021) to facilitate feature interaction. Theoretically, certain parts of the textual and visual features should convey the same information. However, differences in feature spaces make direct mapping between text and image features challenging, requiring an intermediate medium to facilitate the process. Chen et al. (2021) employed a $N\times D$ metric for this purpose, where $N$ refers to the number of memory vectors, and $D$ refers to the dimension of each memory vector. First, the metric was initialized randomly. Image features, text features from generated tokens, and memory metric features then underwent linear transformation to align their feature spaces. Subsequently, distances between the image features and memory metric features, as well as text features and memory metric features, were calculated. The top $K$ metric features with the closest distances to the image or text were selected, respectively. These selected features were then weighted based on these distances and were fed back into an encoder-decoder structure. This metric, serving as a medium, ensured that the selected visual and textual features belong to the same feature space (i.e., memory metric), loosely aligning the visual and textual features and reducing the problems arising from distinct feature spaces. Additionally, the authors tested separate textual and visual memory metrics, which also improved performance. This suggests that regularity patterns may exist within radiology images and reports, and using the memory metric can better capture these patterns to enhance model performance. This point is further supported by Yang et al. (2023), who introduce a memory metric that dynamically captures knowledge during training. This metric is then used as a knowledge base during inference to address the limitation of manually constructing knowledge graphs, as discussed in Section 3.3.1. The memory metric is updated at each step by incorporating textual and visual features, as well as metrics from the previous step, using an attention mechanism.
@@
Two studies (Qin and Song, 2022; You et al., 2022) followed Chen et al. (2021). Wang et al. (2022a) modified it in two ways: 1) they initialized the memory metric by visual and textual features; 2) the cross-modal interaction occurred only among cases with the same label. These two modifications resulted in a notable improvement. Li et al. (2023e) contended that the methods mentioned lack explicit constraints for cross-modal alignments. They considered orthonormal bases as the memory metric and input them along with visual or textual features, into multi-head attention modules. Then the outputs of attention modules were processed by a self-defined gate mechanism. A triplet matching loss was utilized to align the processed visual and textual features. This method slightly improved the results.
Overall, the multimodal fusion and interaction methods reviewed primarily focus on addressing challenges associated with distinct feature spaces. To balance the contributions of different modalities, a common strategy involves automatically learning the weights of contributions during training. However, regarding the other two challenges, many approaches rely heavily on similarity calculations between inputs to identify informative features. This reliance limits their ability to capture complementary information that is critical for maximizing the strengths of each modality. Furthermore, there is often a lack of specialized mechanisms to effectively address various types of noise.
@@ 3.5 Report generation
The last step is report generation, which utilizes extracted features from earlier steps to produce the final reports. The generation methods mainly include decoder-based techniques (Section 3.5.1), retrieval-based techniques (Section 3.5.2), and template-based techniques (Section 3.5.2).
3.5.1 Decoder-based techniques for report generation
A typical method for training decoder-based report generation is autoregressive training. In this process, image (or fused) features and generated tokens (or real text preceding the token awaiting generation) are input into the model to compute the probability distribution for the next token. New tokens are generated by sampling from this distribution, and a loss function is applied between the output distribution and the ground truth. Two main challenges arise in this process: first, how to mitigate information forgetting and extract useful information from long texts to generate new tokens. Second, designing an effective loss function is crucial for guiding text generation and avoiding biases, such as favouring high-frequency words over important medical keywords. The approaches to these two challenges are primarily discussed in Sections 3.5.1 and 3.5.1. Additionally, this section presents two other training strategies: reinforcement learning (Section 3.5.1) and curriculum learning (Section 3.5.1).
Architectures for decoder-based techniques
The decoder decodes the extracted representation of inputs and generates a descriptive report. The mainstream architectures include LSTM (Hochreiter and Schmidhuber, 1997) and Transformer. Compared to LSTM, the Transformer processes the entire sequence simultaneously rather than sequentially. Therefore, the Transformer allows for more efficient parallelization during training and can capture long-range dependencies. In the 100 reviewed papers, the Transformer tends to replace LSTM. Sixty-six works utilized the Transformer as a decoder and only 23 of them utilized LSTMs (12 works) or hierarchical LSTMs (11 works). For the papers published in 2023 and 2024, all encoder-decoder structures used the Transformer as their decoders. There are two ways to modify the decoder and improve model performance by more effectively capturing the information within the text: shortcut connections and memory-driven transformer.
@@ 3.5.1 Decoder-based techniques for report generation
Memory-driven Transformer: We would like to highlight the Memory-driven Transformer (R2Gen) proposed by Chen et al. (2020). It has been increasingly popular in recent years. The R2Gen introduces a memory module and a memory-driven conditional layer normalization module into the Transformer decoder architecture. The design of the memory module hypothesizes that diverse images exhibit similar patterns in their radiological reports, thereby serving as valuable references for each other. Building a memory matrix can capture this pattern and transfer it during training. Specifically, similar to that in Section 3.3.1, a matrix is randomly initialized and is updated using the gate mechanism based on the matrix from the last step and generated reports. The layer normalization is designed to integrate the outputs of the memory module into the decoder.
@@
Loss functions for decoder-based techniques
The application of an auxiliary loss function can provide additional supervision signals, further enhancing model performance. Two works Wang et al. (2021); Li et al. (2022a) applied an additional constraint between features extracted from the generated and real reports. Wu et al. (2023b) applied a penalty to the predicted probabilities of frequent tokens. Zhang et al. (2023b) created two different versions of an input image by adding noise and feeding them into two networks. An auxiliary loss function ensures the consistency of the outputs produced by the two generators. Wang et al. (2024a) obtained two discriminative regions in an image from the generated words and visual classifier separately, then enforced the consistency between them.
Reinforcement learning for decoder-based techniques
Reinforcement learning offers a method to update the model parameters based on non-differentiable reward functions (Messina et al., 2022). In report generation, the model is typically viewed as an agent interacting with a generation environment. Input features and previously generated tokens are considered as the states of the environment. The parameters of the model define a policy for generating actions (i.e., next token generation) based on the states, with a reward computed at each step based on the action (Qin and Song, 2022). The model is trained to maximize the reward using various strategies such as the policy gradients-based REINFORCE algorithm (Williams, 1992). This reward-based training is often combined with loss-based approaches, either by alternating between them (Kaur and Mittal, 2022b) or first training with loss, followed by rewards (Qin and Song, 2022). Improvement in the specified evaluation metric (see Section 7) can be considered as the basis for designing rewards, such as the Consensus-based Image Description Evaluation (CIDEr) (Kaur and Mittal, 2022b), the Bilingual Evaluation Understudy (BLEU) (Qin and Song, 2022; Gu et al., 2024), the Metric for Evaluation of Translation with Explicit ORdering (METEOR) (Qin and Song, 2022), the Recall-Oriented Understudy for Gisting Evaluation (ROUGE) (Qin and Song, 2022), BERTScore (Miura et al., 2021), F1 score (Miura et al., 2021; Wu et al., 2023b), and accuracy (Hou et al., 2021b). In addition, Hou et al. (2021b) trained a language fluency discriminator using the ground truth and generated reports, and then utilized the discriminator to provide rewards.
Curriculum learning for decoder-based techniques
Data-level curriculum learning is a strategy that organizes training samples in the order of difficulties, allowing the model to learn in a progressive manner (Soviany et al., 2022). Unlike traditional methods, where training data is shuffled and indiscriminately fed into the model, this approach typically involves: (1) ranking the training data based on a predefined strategy; (2) starting the training process with simpler samples and progressively introducing more complex ones; and (3) training the model on the full dataset to reinforce its learning. The key element of this approach is data ranking (i.e., the first step), which mimics the human learning progression. This prevents the model from being overwhelmed by complex samples early in training, thereby reducing the risk of getting stuck in suboptimal solutions. Moreover, the strategy is applicable to both loss-based and reward-based training methods. Liu et al. (2021a) introduced curriculum learning to the field of report generation by designing specific metrics to evaluate the difficulty of image-report training data. They proposed four metrics named image heuristics, image confidence, text heuristics, and text confidence. Image heuristic evaluates the similarity between input images and normal images. Image confidence indicates the confidence of a classification model. The report heuristic was related to the number of abnormal sentences, and report confidence is evaluated by the negative log-likelihood loss (Xu et al., 2020). Instead of directly summing the four metrics, they adaptively select the most suitable metric using dynamic feedback for each training step, and use the selected metric to choose an appropriate training batch. Experimental results show that incorporating this training strategy into various methods leads to improved outcomes.
In conclusion, the Transformer has replaced LSTM as the dominant architecture. Structural enhancements, such as short connections, improve information flow and alleviate gradient vanishing, though they also increase model complexity. The memory matrix allows cross-referencing between cases, but attention must be paid to memory size, as an excessively large memory can introduce redundancy and degrade performance. Another improvement lies in the loss function. It involves either assigning different weights to words or incorporating auxiliary loss functions; however, these additional supervision signals provide limited information compared to those used in image feature extraction. Reinforcement learning offers a method for optimizing non-differentiable objectives, but focusing on specific rewards (e.g., ROUGE-L) may negatively impact performance on other metrics, making reward function design critical. Additionally, curriculum learning presents a promising approach to organizing training data.
3.5.2 Retrieval-based and template-based techniques for report generation